40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2616     Entry time: Fri Feb 19 10:18:19 2010
Author: Jenne 
Type: Update 
Category: VAC 
Subject: The P1 vac pressure is almost to 3mTorr 

The Vac pressure measured at P1 is at 2.5mTorr.  I expect we'll hit 3mTorr sometime this afternoon, at which point (according to Steve) the interlock will shut the shutter, and we won't have light in the IFO.  Anything which needs to happen with light in the IFO before Monday needs to happen fairly soon.

Attachment 1: VacPressureAlmostShutoffLaser_19Feb2010.png  10 kB  | Show | Show all
ELOG V3.1.3-